OpenStudy (anonymous):

What is the exact value of the square root of the quantity 98 over 12.?

OpenStudy (anonymous):

\[\sqrt{\frac{ 98 }{ 12 }}\]

OpenStudy (anonymous):

I can get to \[\frac{ 7\sqrt{2} }{ 2\sqrt{3} }\] but I get stuck there